OpenStudy (ebydaqueen19):

HELP !! HELP !! It has been found that the distance from a person's eye to the end of the fingers of his outstretched hand is proportional to the distance between his eyes at a 10:1 ratio. If the distance between your eyes is 2.3 inches, what should the distance from your eye to your outstretched fingers be?

OpenStudy (anonymous):

eyes/fingers = 1/10 fingers/eyes = 10/1 fingers = (10)(eyes)

OpenStudy (ebydaqueen19):

THANK YOU VERY MUCH